# Floating-Base Deep Lagrangian Networks

### v1

@@ Abstract
Grey-box methods for system identification combine deep learning with physics-informed constraints, capturing complex dependencies while improving out-of-distribution generalization. Yet, despite the growing importance of floating-base systems such as humanoids and quadrupeds, current grey-box models ignore their specific physical constraints. For instance, the inertia matrix is not only positive definite but also exhibits branch-induced sparsity and input independence. Moreover, the 6×6 composite spatial inertia of the floating base inherits properties of single-rigid-body inertia matrices. As we show, this includes the triangle inequality on the eigenvalues of the composite rotational inertia. To address the lack of physical consistency in deep learning models of floating-base systems, we introduce a parameterization of inertia matrices that satisfies all these constraints. Inspired by Deep Lagrangian Networks (DeLaN), we train neural networks to predict physically plausible inertia matrices that minimize inverse dynamics error under Lagrangian mechanics. For evaluation, we collected and released a dataset on multiple quadrupeds and humanoids. In these experiments, our Floating-Base Deep Lagrangian Networks (FeLaN) achieve highly competitive performance on both simulated and real robots, while providing greater physical interpretability.
@@ I INTRODUCTION
Accurate dynamic models are essential for simulation, control, and state estimation in robotics. Classical system identification (SysID) constructs a white-box model based on physical principles and estimates its parameters from data. This approach requires prior knowledge of the system, but generalizes well as long as the model assumptions hold. In contrast, black-box methods approximate the dynamics directly from data without any prior knowledge. Although flexible, they often generalize poorly outside the training domain and lack interpretability. Grey-box methods combine data-driven models with physical structure. For example, Hamiltonian Neural Networks (HNN) [1] and Deep Lagrangian Networks (DeLaN) [2] embed Hamiltonian and Lagrangian mechanics into deep learning architectures, respectively. By enforcing physical properties such as conservation laws, these formulations improve interpretability, generalizability, and sample efficiency, and have demonstrated success on fixed-base [3] and single rigid-body [4, 5] robots. In addition, learning function approximations allows to learn a physically consistent model when the true system states are partially or entirely unobserved [6], or combine the true states with additional representations for context adaptation [7].
However, naively applying these methods to floating-base systems, such as quadrupeds and humanoids, neglects several physical constraints which may limit the generalizability of the learned model. First of all, such robots contain multiple kinematic chains, leading to an inertia matrix with a sparse block structure where certain blocks are independent of specific joint inputs. Furthermore, the composite spatial inertia matrix [8], i.e., the part of the inertia matrix that maps floating-base accelerations to the total wrench applied at the base, should respect a triangular inequality [9]. Even for DeLaN, which directly predicts the Cholesky decomposition of the inertia matrix, imposing these additional constraints is challenging: sparsity and input-independencies are not preserved in the Cholesky matrix [10] and enforcing the triangular inequality is non-trivial. Hence, grey-box models for floating-base systems remain limited to simplified dynamics [11, 12].
In this work, we introduce a novel parameterization of the inertia matrix of floating-base system based on the reordered Cholesky factorization [10]. Our parameterization ensures positive-definiteness; branch-induced sparsity and input-independencies; the triangular inequality of the composite spatial inertia matrix; and the stationarity of the system total mass. Based on DeLaN, we propose the Floating-Base Deep Lagrangian Networks (FeLaN), a novel grey-box method for the identification of floating-base system dynamics. We validate our methods in several simulation environments and on multiple real quadrupeds and humanoids, namely on Unitree Go2 [13], Boston Dynamics Spot [14], Pal Robotics TALOS [15], and HyQReal2, the updated version of HyQReal1 [16].
@@ II Related Work
Classical SysID techniques for rigid-body systems usually estimate the kinematic parameters (e.g., link lengths, joint axes) separately from the dynamic parameters (e.g., mass, center of mass, inertia), either by prior estimation or by assuming they are given [17]. This separation is motivated by the fact that the equations of motion are linear in the inertial parameters, which enables estimation via linear least squares assuming persistent excitation [18]. An alternative approach is to employ differentiable simulators to directly identify all system parameters through stochastic gradient descent [19, 20]. For floating-base systems, different sensor modalities were explored, such as joint torques [21], joint torques and force sensors [22], or force plates [23].
@@
The integration of physics into deep learning demonstrated a useful inductive bias, especially in small data regimes [28]. For floating-base robots, [11] used DeLaN to estimate simplified formulations of a quadruped’s inertia matrix for model-based locomotion. However, only the positive definiteness of the inertia matrix is enforced, which is a necessary but not sufficient condition for full physical consistency. The composite spatial inertia of a floating-base system must also satisfy the triangle inequality for the principal moments of the rotational inertia, as in the single rigid body case presented in [9].
@@
Floating-base systems may have multiple kinematic branches, which are only indirectly coupled through the base, resulting in a partial functional decoupling. In an attempt to exploit this property, [34] learns an inverse dynamics function for each robot leg. However, they included contact state variables and a projection of the base position in the leg frame, which requires information regarding the robot kinematics. Furthermore, the partial functional decoupling is truly described by the branch-induced sparsity of the inertia matrix [8], which was first exploited by [10] to efficiently factorize a given inertia matrix.
Unlike structural constraints, Sparse Identification of Nonlinear Dynamics (SINDy) [35] learns parsimonious models through optimization, leveraging the inherent sparsity of physical laws. For example, [12] combines SINDy with priors on quadruped jumping dynamics to estimate a low-dimensional model equivalent to a single rigid body. While this enforces sparsity, it fails to capture branch-induced sparsity and to guarantee the properties of the composite spatial inertia.
@@ III-A Notation and Definitions
Scalar and scalar-value functions are denoted by lowercase letters; vector and vector-valued are denoted with bold lowercase letters; and matrices with uppercase letters. The $\mathrm{n}\times\mathrm{n}$ identity matrix is denoted as $\mathbf{1}_{\mathrm{n}}$. $\mathbf{S}(\mathbf{a})$ and $\mathbf{S}_{\mathbf{a}}$ denote the skew-symmetric operator over the vector $\mathbf{a}$. $\mathbf{A}\succeq 0$ and $\mathbf{A}\succ 0$ represents that the symmetric matrix $\mathbf{A}$ is positive semidefinite and definite, respectively. The superscript W denotes a variable in the world frame, whereas all other variables are expressed in the base frame. The dependence of matrices and vectors on configuration variables is written explicitly at first mention. For brevity, these dependencies are omitted in subsequent expressions but remain implicitly valid.
@@ III-B Lagrangian Mechanics
Lagrangian mechanics describes the dynamics of mechanical systems by their energy. The Lagrangian $\mathcal{L}$ is defined as the difference between the kinetic $K$ and potential energy $P$. For rigid body systems, the Lagrangian takes the form
@@
where $\boldsymbol{\nu}$ and $\dot{\boldsymbol{\nu}}$ are the generalized positions and velocities, respectively, and $\mathbf{H}(\boldsymbol{\nu})$ is the positive definite inertia matrix.
@@ III-C Deep Lagrangian Networks
Based on (1), DeLaN [36] approximates the joint space inertia matrix $\mathbf{H}$ and $P$ using two networks to estimate the equations of motion (3) of a fixed-base robot. To guarantee physical consistency, i.e., $\mathbf{H}\succ 0$, DeLaN parameterizes the inertia matrix in terms of its Cholesky decomposition, i.e.
@@ IV Physically Consistent Inertia Matrix Parametrization
In this section, we first investigate the properties of the $\mathbf{H}$ of a floating-base system, and extend the concept of fully physical consistency to the composite spatial inertia. We then propose an unconstrained parametrization of $\mathbf{H}$ that address all the investigated physical constraints.
@@ IV-A Inertia Matrix of a Floating-Base system
For floating-base systems, the generalized coordinates account for both the base pose $\mathbf{q}_{\mathbf{B}}\in\mathrm{SE}(3)$, and the joint position $\mathbf{q}\in\mathbb{R}^{\mathrm{n}_{\mathrm{q}}}$, leading to $\boldsymbol{\nu}=[\mathbf{q}_{\mathbf{B}};\mathbf{q}]$. The top-left $6\times 6$ block matrix of $\mathbf{H}$ corresponds to the composite spatial inertia $\mathbf{H}_{\mathbf{B}}(\mathbf{q})$ of the whole floating-base system [8]. Equivalent to (5), the total mass $\mathrm{m}\in\mathbb{R}$, the total first mass moment $\mathbf{h}(\mathbf{q})\in\mathbb{R}^{3}$, and the composite rotational inertia $\mathbf{I}_{\mathbf{B}}(\mathbf{q})\in\mathbb{R}^{3\times 3}$, define $\mathbf{H}_{\mathbf{B}}(\mathbf{q})$ in the base frame as
@@
Unlike the spatial inertia of a single rigid body, both $\mathbf{h}(\mathbf{q})$ and $\mathbf{I}_{\mathbf{B}}(\mathbf{q})$ depend on the joint configuration $\mathbf{q}$ due the transformation from body to base frame, whereas $\mathrm{m}$ remains constant and results in an isotropic linear inertia. Since $\mathbf{H}_{\mathbf{B}}(\mathbf{q})$ has the same structure as the spatial inertia of a single rigid body (5), $\mathbf{H}_{\mathbf{B}}(\mathbf{q})\succ 0$ is a necessary but not sufficient condition for full physical consistency, as established by Lemma 1.
@@
A composite inertia matrix is fully physical consistent only if $\mathrm{m}>0$ and $\mathbf{I}_{\mathbf{B}}(\mathbf{q})$ satisfies the triangle inequality.
@@
$\mathbf{I}_{i}(\mathbf{q})=\mathbf{R}_{i}(\mathbf{q})\mathbf{I}_{b}{}_{i}\mathbf{R}_{i}(\mathbf{q})+\mathrm{m}_{i}\mathbf{S}(\mathbf{r}_{i}(\mathbf{q}))\mathbf{S}(\mathbf{r}_{i}(\mathbf{q}))\text{,}$ (10)
where $\mathbf{R}_{i}(\mathbf{q})\in\mathrm{SO}(3)$ is the rotation from body and base frame, $\mathbf{r}_{i}$ is the body position w.r.t. the base. The parallel axis theorem preserves the triangular inequality, as it corresponds to a change of reference frame [9]. From (7), the trace of the composite rotational inertia satisfies a minimum bound:
@@
Note that by expressing $\mathbf{H}_{\mathbf{B}}(\mathbf{q})$ in the base frame, the inertia matrix $\mathbf{H}(\mathbf{q})$ depends only on the joint positions $\mathbf{q}$, but not on the entire generalized coordinates $\boldsymbol{\nu}$, resulting in translational and rotational invariances regarding the base.
Due to the branched-induced sparsity, the joint inertia $\mathbf{H}_{k}$ of the $k$th kinematic branch is a function only of the joints of the $k$th branch, along with the linear and rotational inertial coupling with the robot’s base.
@@ IV-B Branch-Induced Sparsity
For fixed-base robots with a single kinematic branch, a positive definite inertia matrix can be estimated from the Cholesky factor (4[2]. However, as established by Lemma 1, $\mathbf{H}\succ 0$ is a necessary but not a sufficient condition for a physically consistent $\mathbf{H}$ of a floating-base system.
The main limitation is that the standard Cholesky factorization (4) always results in a dense factor $\mathbf{C}(\mathbf{q})$ [10], not holding the branched-induced sparsity and not necessarily the structure of $\mathbf{H}_{\mathbf{B}}(\mathbf{q})$.
To exploit the branch-induced sparsity and efficiently factorize a given $\mathbf{H}$, [10] proposes to use a reordered Cholesky factorization
@@
where $\mathbf{L}_{\mathrm{L}}(\mathbf{q})$, $\mathbf{L}_{\mathrm{R}\mathrm{L}}(\mathbf{q})$, $\mathbf{L}_{\mathrm{R}}(\mathbf{q})$ are a function of all joints $\mathbf{q}$; $\mathbf{L}_{k\mathrm{L}}(\mathbf{q}_{k})$, $\mathbf{L}_{k\mathrm{R}}(\mathbf{q}_{k})$, $\mathbf{L}_{k}(\mathbf{q}_{k})$ depend only of $\mathbf{q}_{k}$ of the $k$th kinematic branch. For brevity, we omit explicitly writing the dependence of matrices and vectors on the joint positions $\mathbf{q}$ in subsequent expressions; this dependence remains valid.
@@ IV-D Inertia Triangle Inequality
where $\mathbf{W}(\mathbf{q})=[\mathbf{L}_{1\mathrm{R}}(\mathbf{q}_{1}),\dots,\mathbf{L}_{{\mathrm{n}_{\mathrm{k}}}\mathrm{R}}(\mathbf{q}_{\mathrm{n}_{\mathrm{k}}})]$. Hence, we can express $\mathbf{L}_{\mathrm{R}}$ in terms of $\boldsymbol{\Sigma}_{\mathrm{R}}$ and $\mathbf{W}(\mathbf{q})$ as
@@
$\mathbf{L}_{\mathrm{R}}$ $=\mathrm{reorderedChol}(\mathbf{D})\text{ with }\mathbf{D}\succ 0\text{.}$ (19)
The factorization algorithm $\mathrm{reorderedChol}$ is proposed by [10] that factorizes an inertia matrix using the kinematic chain information to exploit the branch-induced sparsity.

### v2

@@ Abstract
Grey-box methods for system identification combine deep learning with physics-informed constraints, capturing complex dependencies while improving out-of-distribution generalization. Despite the growing importance of floating-base systems such as humanoids and quadrupeds, current grey-box models ignore their specific physical constraints. For instance, the inertia matrix is not only positive definite but also exhibits branch-induced sparsity and input independence. Moreover, the 6×6 composite spatial inertia of the floating base inherits properties of single-rigid-body inertia matrices. As we show, this includes the triangle inequality on the eigenvalues of the composite rotational inertia. To address the lack of physical consistency in deep learning models of floating-base systems, we introduce a parameterization of inertia matrices that satisfies all these constraints. Inspired by Deep Lagrangian Networks (DeLaN), we train neural networks to predict physically plausible inertia matrices that minimize inverse dynamics error under Lagrangian mechanics. For evaluation, we collected and released a dataset on multiple quadrupeds and humanoids. In these experiments, our Floating-Base Deep Lagrangian Networks (FeLaN) achieve better overall performance on both simulated and real robots, while providing greater physical interpretability.
@@ I INTRODUCTION
Accurate dynamic models are essential for simulation, control, and state estimation in robotics. Classical system identification (SysID) constructs a white-box model based on physical principles and estimates its parameters from data. This approach requires prior knowledge of the system, but generalizes well as long as the model assumptions hold. In contrast, black-box methods approximate the dynamics directly from data without any prior knowledge. Although flexible, they often generalize poorly outside the training domain and lack interpretability. Grey-box methods combine data-driven models with physical structure. For example, Hamiltonian Neural Networks (HNN) [16] and Deep Lagrangian Networks (DeLaN) [24] embed Hamiltonian and Lagrangian mechanics into deep learning architectures, respectively. By enforcing physical properties such as conservation laws, these formulations improve interpretability, generalization, and sample efficiency, and have demonstrated success on fixed-base [23] and single rigid-body [10, 1] robots. Moreover, learned function approximators can yield physically consistent models even with partially or fully unobserved states [36], and can be combined with latent representations [31].
Figure 1: Floating-Base Deep Lagrangian Networks (FeLaN) for Unitree Go2 quadruped.
However, naively applying these methods to floating-base systems, such as quadrupeds and humanoids, neglects several physical constraints which may limit the generalization of the learned model. These robots have multiple kinematic chains, leading to a sparse inertia matrix with blocks dependent on specific joint inputs. The composite spatial inertia matrix [13], which maps floating-base accelerations to the total wrench applied at the base, must respect a triangle inequality [37]. Even for DeLaN, which predicts the Cholesky decomposition of the inertia matrix, imposing these additional constraints is challenging: sparsity and input-independencies are not preserved in the Cholesky matrix [12] and enforcing the triangle inequality is non-trivial. Hence, grey-box models for floating-base robots remain limited to simplified dynamics [19, 8].
In this work, we introduce a novel parameterization of the inertia matrix of floating-base systems based on the reordered Cholesky factorization [12]. Our parameterization ensures positive definiteness; branch-induced sparsity and input-independencies; the triangle inequality of the composite spatial inertia; and the stationarity of the system total mass. Based on DeLaN, we propose Floating-Base Deep Lagrangian Networks (FeLaN), a novel grey-box method for identifying the dynamics of floating-base systems, depicted in Fig. 1. We validate our method in several simulation environments and on multiple real quadrupeds and humanoids: Unitree Go2 [38], Boston Dynamics Spot [5], Pal Robotics Talos [33], and HyQReal2, the updated version of HyQReal [32].
@@ II Related Work
Classical SysID techniques for rigid-body systems usually estimate the kinematic parameters (e.g., link lengths, joint axes) separately from the dynamic parameters (e.g., mass, center of mass, inertia), either by prior estimation or by assuming they are given [22]. This separation is motivated by the linearity of the equations of motion in the inertial parameters, enabling estimation via linear least squares under persistent excitation [21]. An alternative approach is to employ differentiable simulators to directly identify all system parameters through stochastic gradient descent [25, 34]. For floating-base systems, prior work has explored different sensing modalities, including joint torques [18], joint torques with force sensors [28], and force plates [4].
@@
Integrating physics into deep learning demonstrated a useful inductive bias, especially in small data regimes [39]. For floating-base robots, [19] used DeLaN to estimate a simplified formulation of a quadruped’s inertia matrix for model-based locomotion. However, only positive definiteness of the inertia is enforced, a necessary but not sufficient condition for full physical consistency, as we prove in Sec. IV-A.
@@
Floating-base systems may have multiple kinematic branches only indirectly coupled through the base, yielding partial functional decoupling. To exploit this property, [20] learns an inverse dynamics function for each leg, including the base position expressed in the leg frame, which requires kinematic information. In fact, the partial functional decoupling is truly described by the branch-induced sparsity of the inertia matrix [13], which was first leveraged by [12] for efficient inertia matrix factorization.
Unlike structural constraints, Sparse Identification of Nonlinear Dynamics (SINDy) [7] promotes sparsity through optimization. [8] combines linear autoencoders with SINDy to learn a reduced-order quadruped model. While this enforces sparsity, it fails to capture branch-induced sparsity and to guarantee the composite spatial inertia properties.
@@ III-A Notation and Definitions
The $\mathrm{n}\times\mathrm{n}$ identity matrix is denoted as $\mathbf{1}_{\mathrm{n}}$. $\mathbf{S}(\mathbf{a})$ and $\mathbf{S}_{\mathbf{a}}$ denote the skew-symmetric operator over the vector $\mathbf{a}$. $\mathbf{A}\succeq 0$ and $\mathbf{A}\succ 0$ represents that the symmetric matrix $\mathbf{A}$ is positive semidefinite and definite, respectively. The superscript W denotes a variable in the world frame, whereas all other variables are expressed in the base frame. The dependence of matrices and vectors on configuration variables is written explicitly at first mention.
@@ III-B Lagrangian Mechanics
Lagrangian mechanics formulates rigid-body dynamics in terms of the Lagrangian
@@
where $\boldsymbol{\nu}$ and $\dot{\boldsymbol{\nu}}$ are generalized positions and velocities, $K$ and $P$ are the kinetic and potential energies, and $\mathbf{H}(\boldsymbol{\nu})$ is the positive-definite inertia matrix.
@@ III-C Deep Lagrangian Networks
Based on (1), DeLaN [26] approximates the joint-space inertia matrix $\mathbf{H}$ and $P$ using two networks to estimate the equations of motion (3) of a fixed-base robot. To guarantee physical consistency, i.e., $\mathbf{H}\succ 0$, DeLaN parameterizes the inertia matrix in terms of its Cholesky decomposition, i.e.
@@ IV Physically Consistent Inertia Matrix Parametrization
In this section, we investigate the floating-base inertia matrix $\mathbf{H}$ and extend full physical consistency to the composite spatial inertia. We then propose an unconstrained parametrization of $\mathbf{H}$ that addresses all identified constraints.
@@ IV-A Inertia Matrix of a Floating-Base system
For floating-base systems, the generalized coordinates account for both base pose $\mathbf{q}_{\mathbf{B}}\in\mathrm{SE}(3)$, and joint position $\mathbf{q}\in\mathbb{R}^{\mathrm{n}_{\mathrm{q}}}$, leading to $\boldsymbol{\nu}=[\mathbf{q}_{\mathbf{B}};\mathbf{q}]$. The top-left $6\times 6$ block matrix of $\mathbf{H}$ corresponds to the composite spatial inertia $\mathbf{H}_{\mathbf{B}}(\mathbf{q})$ of the whole system [13]. As in (5), $\mathbf{H}_{\mathbf{B}}$ is defined in the base frame from the total mass $\mathrm{m}$, total first mass moment $\mathbf{h}(\mathbf{q})$, and composite rotational inertia $\mathbf{I}_{\mathbf{B}}(\mathbf{q})$ as
@@
Unlike the spatial inertia of a single rigid body, both $\mathbf{h}(\mathbf{q})$ and $\mathbf{I}_{\mathbf{B}}(\mathbf{q})$ depend on the joint configuration $\mathbf{q}$ due to the transformation from body to base frame, whereas $\mathrm{m}$ remains constant and results in an isotropic linear inertia. Since $\mathbf{H}_{\mathbf{B}}$ has the same structure as the spatial inertia of a single rigid body (5), $\mathbf{H}_{\mathbf{B}}\succ 0$ is a necessary but not sufficient condition for full physical consistency, as established by Lemma 1.
@@
A composite inertia matrix is fully physically consistent only if $\mathrm{m}>0$ and $\mathbf{I}_{\mathbf{B}}(\mathbf{q})$ satisfies the triangle inequality.
@@
$\mathbf{I}_{i}(\mathbf{q})=\mathbf{R}_{i}(\mathbf{q})\mathbf{I}_{b}{}_{i}\mathbf{R}_{i}(\mathbf{q})^{\mathrm{\scriptscriptstyle T}}+\mathrm{m}_{i}\mathbf{S}(\mathbf{r}_{i}(\mathbf{q}))\mathbf{S}(\mathbf{r}_{i}(\mathbf{q}))^{\mathrm{\scriptscriptstyle T}}\text{,}$ (10)
where $\mathbf{R}_{i}(\mathbf{q})\in\mathrm{SO}(3)$ is the rotation from body to base frame, $\mathbf{r}_{i}$ is the body position w.r.t. the base. The parallel axis theorem preserves the triangle inequality, as it corresponds to a change of reference frame [37]. From (7), the trace of the composite rotational inertia satisfies a minimum bound:
@@
Note that by expressing $\mathbf{H}_{\mathbf{B}}$ in the base frame, $\mathbf{H}$ depends only on $\mathbf{q}$ rather than on $\boldsymbol{\nu}$, yielding translational and rotational invariance with respect to the base coordinates.
Due to the branch-induced sparsity, the joint inertia $\mathbf{H}_{k}$ of the $k$th kinematic branch is a function only of the $k$th branch joints $\mathbf{q}_{k}$, along with the linear and rotational inertial coupling with the robot’s base.
@@ IV-B Branch-Induced Sparsity
For fixed-base robots with a single kinematic branch, a positive-definite inertia matrix can be estimated from the Cholesky factor (4[24]. However, as established by Lemma 1, $\mathbf{H}\succ 0$ is a necessary but not a sufficient condition for a physically consistent $\mathbf{H}$ of a floating-base system.
The main limitation is that the standard Cholesky factorization (4) typically results in a dense factor $\mathbf{C}(\mathbf{q})$ [12], not holding the branched-induced sparsity and not necessarily the structure of $\mathbf{H}_{\mathbf{B}}(\mathbf{q})$. To exploit the branch-induced sparsity and efficiently factorize a given $\mathbf{H}$, [12] proposes to use a reordered Cholesky factorization
@@
where $\mathbf{L}_{\mathrm{L}}(\mathbf{q})$, $\mathbf{L}_{\mathrm{R}\mathrm{L}}(\mathbf{q})$, $\mathbf{L}_{\mathrm{R}}(\mathbf{q})$ are a function of all joints $\mathbf{q}$; $\mathbf{L}_{k\mathrm{L}}(\mathbf{q}_{k})$, $\mathbf{L}_{k\mathrm{R}}(\mathbf{q}_{k})$, $\mathbf{L}_{k}(\mathbf{q}_{k})$ depend only on $\mathbf{q}_{k}$ of the $k$th kinematic branch.
@@ IV-D Triangle Inequality
where $\mathbf{W}(\mathbf{q})=[\mathbf{L}_{1\mathrm{R}}(\mathbf{q}_{1});\dots;\mathbf{L}_{{\mathrm{n}_{\mathrm{k}}}\mathrm{R}}(\mathbf{q}_{\mathrm{n}_{\mathrm{k}}})]$. Hence, we can express $\mathbf{L}_{\mathrm{R}}$ in terms of $\boldsymbol{\Sigma}_{\mathrm{R}}$ and $\mathbf{W}(\mathbf{q})$ as
@@
$\mathbf{L}_{\mathrm{R}}$ $=\mathrm{reorderedChol}(\mathbf{D})\text{ with }\mathbf{D}\succ 0\text{,}$ (19)
where $\mathrm{reorderedChol}$ is the factorization proposed by [12].